# Stochastic trade-offs and the emergence of diversification in E. coli evolution experiments

### v2

@@ V Acknowledgements
This work has been supported by grant PID2020-113681GB-I00 of the Spanish Ministry and Agencia Estatal de Investigación (MCIN/AEI/10.13039/501100011033). R.C.L. acknowledges funding from the Spanish Ministry and AEI, under Grant No. FPU19/03887 and from the Arqus alliance of european universities program 6.5. SS and SA acknowledge the support of the NBFC to the University of Padova, funded by the Italian Ministry of University and Research, PNRR, Missione 4 Componente 2, “Dalla ricerca all’impresa”, Investimento 1.4, Project CN00000033. The authors warmly thank Amos Maritan for insightful discussions and for contributions to the early stages of this research.
VI Methods
@@ VI.1 Branching feasibility plots for monomorphic populations
Indeed, the phenotype x evolves according to the canonical equation (Eq.(5)), i.e. it attempts to climb the local selection gradient $d\textbf{x}/dt\propto\textbf{g}(\textbf{x})$ [26]. Nevertheless, the presence of a trade-off between the evolving phenotypes, that can be written in general as $H(\textbf{V})=0$, constrains the evolution to the projection of the gradient onto the trade-off curve [30]:
@@
In the two-dimensional case under consideration, one has that $\textbf{x}=(x_{1},x_{2})$ and $x_{2}=h(x_{1})$ where $h$ is an arbitrary trade-off function. If evolution is constrained to follow a hard trade-off curve, it reaches an endpoint (singular point) if at some point the selection gradient is perpendicular to the tangent vector of the trade-off function, h, i.e. $\textbf{g}\cdot\textbf{h}=0$ (see points A and B in 5a). Therefore, one can construct curves that, at any point in phenotypic space, are orthogonal to the local selection gradient and hence to all resultant evolutionary trajectories. These curves (blue lines in 5a), called A-boundaries, are then independent from the trade-off and depend exclusively on the characteristics of the model. Is only when we are to find the specific fixed point(s) of the evolutionary dynamics that we have to specify the trade-off function. Singular points can then be seen as those in which the slopes of the A boundary, $S_{A}$, and the trade-off function, $S_{T}$, are equal or equivalently when both curves are tangent to each other. Besides, A boundaries help to distinguish in a simple geometrical way whether the fixed point is convergence stable, and as such no neighbouring phenotypes can be reached by directional evolution, (A point in Fig.5a) or unstable, where neighbouring phenotypes can indeed be reached by evolution, (B point in Fig.5a).
@@
Thus, the relative local curvatures of A-boundaries, I-boundaries, and trade-off curves at singular points determine the different possible evolutionary outcomes. In order to compare the curvatures, a convention is needed. Since, by definition, the local selection gradient at a singular point is perpendicular to the local I-boundary, A-boundary, and trade-off curve: we define positive curvature when looking along the local selection gradient the curve is convex, otherwise the curvature is negative. Denoting the curvatures of the A, I boundaries and the trade-off by $C_{A}$, $C_{I}$ and $C_{T}$ respectively we can then conclude that a singular phenotype is convergence table (unstable) if $C_{T}<C_{A}$ ($C_{T}>C_{A}$) and locally evolutionary stable (unstable) if $C_{T}<C_{I}$ ($C_{T}>C_{I}$). Thus, the continuously convergence strategy and evolutionary branching points can be easily distinguished (see Fig.5b). Analytical details corresponding to these geometric insights are provided in [44].
To discriminate the conditions for which an initially monomorphic population can become polymorphic. For this the initial point in phenotypic space must converge to an evolutionary branching point which, as we have shown, requires $C_{I}<C_{T}<C_{A}$. One can therefore visualize where in phenotypic space evolutionary branching is more independent from the trade-off by plotting at each point the difference $C_{A}-C_{I}$ (see Fig.3a in the main text).
@@ VI.3 Evolutionary algorithm
Here we follow a slightly modified version of the evolutionary algorithm employed in [54]. To simulate the evolution of the whole population, one needs first to extend the system of ecological dynamics Eqs.(1-3) for an arbitrary number M of sub-populations:
@@ VI.5 Evolutionary branching with a stochastic trade-off
In order to demonstrate that the evolutionary-branching feasibility plot (Fig.3a) does indeed shows where in phenotypic space is more likely to have branching, we showed the percentage of realizations that end up causing branching at the specified points in phenotypic space (Fig.3b). Here we explain how we have obtained Fig.3b: (i) we choose an average curve of the stochastic trade-off that has the right properties to deterministically (in absence of the stochastic trade-off amplitude, that is, $\sigma=0$) originate evolutionary branching at a certain point, $\textbf{V}^{s}=(V^{s}_{g},V^{s}_{a})$ (for example the points marked with a cross in Fig.3a). This can be easily done as detailed in Methods Sec. VI.2. (ii) Then, once we have the stochastic trade-off defined with a certain $\sigma$ and $\alpha$, we initiate the evolutionary algorithm with a monomorphic population whose phenotype, is equal to that of the selected point, $\textbf{V}^{\text{init}}=\textbf{V}^{s}$. We do it in this way to be able to compare the branching times for the different point that we select. (iii) At $t^{\text{max}}_{\text{evo}}$ (which can vary as specified in the label of Fig.3) we determine whether two phenotypically distinguishable groups of subpopulations emerged as described in Methods Sec.VI.4. (iv) We repeat this process for all the different realizations and stochastic trade-off amplitudes $\sigma$.

### v3

@@ V Acknowledgements
This work has been supported by Grant No. PID2020-113681GB-I00 of the Spanish Ministry and Agencia Estatal de Investigación MICIU/AEI/10.13039/501100011033. It has also been supported by Grant No. PID2023-149174NB-I00 financed by MICIU/AEI/10.13039/501100011033 and EDRF funds. R.C.L. acknowledges funding from the Spanish Ministry and AEI, under Grant No. FPU19/03887 and from the Arqus Alliance of European Universities program 6.5. S.A. acknowledges the support of the NBFC to the University of Padova, funded by the Italian Ministry of University and Research, PNRR, Missione 4 Componente 2, “Dalla ricerca all’impresa,” Investimento 1.4, Project No. CN00000033. S.S. acknowledges financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call for Tender No. 104 by the Italian Ministry of University and Research (MUR), funded by the European Union—NextGenerationEU—Project Title: Anchialos: Diversity, function, and resilience of Italian coastal aquifers upon global climatic changes—CUP C53D23003420001 Grant Assignment Decree No. 1015 adopted by the Italian Ministry of Ministry of University and Research (MUR). The authors warmly thank Amos Maritan for insightful discussions and for contributions to the early stages of this research.
VI Supplementary Material
@@ VI.1 Branching feasibility plots for monomorphic populations
Indeed, the phenotype x evolves according to the canonical equation, i.e. it attempts to climb the local selection gradient $d\textbf{x}/dt\propto\textbf{g}(\textbf{x})$ [26]. Nevertheless, the presence of a trade-off between the evolving phenotypes, that can be written in general as $H(\textbf{V})=0$, constrains the evolution to the projection of the gradient onto the trade-off curve [30]:
@@
In the two-dimensional case under consideration, one has that $\textbf{x}=(x_{1},x_{2})$ and $x_{2}=h(x_{1})$ where $h$ is an arbitrary trade-off function. If evolution is constrained to follow a hard trade-off curve, it reaches an endpoint (singular point) if at some point the selection gradient is perpendicular to the tangent vector of the trade-off function, h, i.e. $\textbf{g}\cdot\textbf{h}=0$ (see points A and B in Fig. 5a). Therefore, one can construct curves that, at any point in phenotypic space, are orthogonal to the local selection gradient and hence to all resultant evolutionary trajectories. These curves (blue lines in 5a), called A-boundaries, are then independent from the trade-off and depend exclusively on the characteristics of the model. Is only when we are to find the specific fixed point(s) of the evolutionary dynamics that we have to specify the trade-off function. Singular points can then be seen as those in which the slopes of the A boundary, $S_{A}$, and the trade-off function, $S_{T}$, are equal or equivalently when both curves are tangent to each other. Besides, A boundaries help to distinguish in a simple geometrical way whether the fixed point is convergence stable, and as such no neighbouring phenotypes can be reached by directional evolution, (A point in Fig.5a) or unstable, where neighbouring phenotypes can indeed be reached by evolution, (B point in Fig.5a).
@@
Thus, the relative local curvatures of A-boundaries, I-boundaries, and trade-off curves at singular points determine the different possible evolutionary outcomes. In order to compare the curvatures, a convention is needed. Since, by definition, the local selection gradient at a singular point is perpendicular to the local I-boundary, A-boundary, and trade-off curve: we define positive curvature when looking along the local selection gradient the curve is convex, otherwise the curvature is negative. Denoting the curvatures of the A, I boundaries and the trade-off by $C_{A}$, $C_{I}$ and $C_{T}$ respectively we can then conclude that a singular phenotype is convergence stable (unstable) if $C_{T}<C_{A}$ ($C_{T}>C_{A}$) and locally evolutionary stable (unstable) if $C_{T}<C_{I}$ ($C_{T}>C_{I}$). Thus, the continuously convergence strategy and evolutionary branching points can be easily distinguished (see Fig.5b). Analytical details corresponding to these geometric insights are provided in [44].
In is especially interesting to discriminate the conditions for which an initially monomorphic population can become polymorphic. For this to occur, evolution must converge to an evolutionary branching point which, as we have shown, requires $C_{I}<C_{T}<C_{A}$. One can therefore visualize where in phenotypic space evolutionary branching is more independent from the trade-off curve by plotting at each point the difference $C_{A}-C_{I}$ (see Fig. 3a in the main text).
@@ VI.3 Evolutionary algorithm
Here we follow a slightly modified version of the evolutionary algorithm employed in [54]. To simulate the evolution of the whole population, one needs first to extend the system of ecological dynamics Eqs.(1-3) of the main text for an arbitrary number M of sub-populations:
@@ VI.5 Evolutionary branching with a stochastic trade-off
In order to demonstrate that the evolutionary-branching feasibility plot (Fig. 3a in the main text) does indeed shows where in phenotypic space is more likely to have branching, we showed the percentage of realizations that end up causing branching at the specified points in phenotypic space (Fig. 3b in the main text). Here we explain how we have obtained Fig. 3b in the main text: (i) we choose an average curve of the stochastic trade-off that has the right properties to deterministically (in absence of the stochastic trade-off amplitude, that is, $\sigma=0$) originate evolutionary branching at a certain point, $\textbf{V}^{s}=(V^{s}_{g},V^{s}_{a})$ (for example the points marked with a cross in Fig. 3a in the main text). This can be easily done as detailed in Methods Sec. VI.2. (ii) Then, once we have the stochastic trade-off defined with a certain $\sigma$ and $\alpha$, we initiate the evolutionary algorithm with a monomorphic population whose phenotype, is equal to that of the selected point, $\textbf{V}^{\text{init}}=\textbf{V}^{s}$. We do it in this way to be able to compare the branching times for the different point that we select. (iii) At $t^{\text{max}}_{\text{evo}}$ (which can vary as specified in the label of Fig. 3 in the main text) we determine whether two phenotypically distinguishable groups of subpopulations emerged as described in Methods Sec.VI.4. (iv) We repeat this process for all the different realizations and stochastic trade-off amplitudes $\sigma$.